OpenStudy (anonymous):

HELP NEEDED!!! WILL FAN AND MEDAL!!!! I need help understanding how to prove identities.

OpenStudy (anonymous):

I have a specific example. Please help...

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

\[\tan \frac{ \cos \theta }{ 1+ \sin \theta }=\frac{ 1 }{ \cos \theta }\]

OpenStudy (anonymous):

It was actually supposed to be tangent theta times cosine theta over one plus sine theta

OpenStudy (anonymous):

I have all of the identities written out so I can easily refer to them, but the problem I am having is learning how to combine the like terms.

OpenStudy (anonymous):

\[\frac{ \sin \theta }{ \cos \theta }+\frac{ \cos \theta }{ 1+ \sin \theta }\] this is what I need help combining because I have already converted Tan

OpenStudy (anonymous):

what the heck is the tan doinog there ?

OpenStudy (anonymous):

The tan was there because our teacher was trying to get us to recreate all of the identities to where they equal one another

OpenStudy (anonymous):

no don't convert to tangent in fact, if it was tangent, you would convert to \(\frac{\sin(x)}{\cos(x)}\)

OpenStudy (anonymous):

lets do a tiny bit of algebra first

OpenStudy (anonymous):

That's what I did and that was my farthest point in this.

OpenStudy (anonymous):

ok i got that wrong, hold the phone

OpenStudy (anonymous):

\[\frac{b}{a}+\frac{a}{1+b}\] that's better can you add these?

OpenStudy (anonymous):

I have no idea how to... or at least I can't remember. I know with multiplying you foil it, if you divide you use the kcf format :/ but adding and subtracting makes no sense to me because all I know is a common denominator is required

OpenStudy (anonymous):

lol that is always the problem, the algebra

OpenStudy (anonymous):

here is the one true way to add fractions \[\huge \frac{A}{B}+\frac{C}{D}=\frac{AD+BC}{BD}\]

OpenStudy (anonymous):

o.o"

OpenStudy (anonymous):

in your case you will have \[\frac{b(1+b)+a^2}{a(1+b)}\]

OpenStudy (anonymous):

I am getting more lost by the minute...

OpenStudy (anonymous):

that is the way to add fractions if they are numbers\[\frac{2}{7}+\frac{3}{5}=\frac{2\times 5+7\times 3}{7\times 5}\]

OpenStudy (anonymous):

and it is also the way to add fractions if they have variables it is the only way to do this, by using algebra to add

OpenStudy (anonymous):

ohhhh ok, I get it, sort of.

OpenStudy (anonymous):

forget that "least common denominator" nonsense you were taught that is the way to add, you cannot avoid it

OpenStudy (anonymous):

so ready to start again?

OpenStudy (anonymous):

Yeah I suppose

OpenStudy (anonymous):

using the one true way to add \[\frac{b}{a}+\frac{a}{1+b}\] you get \[\frac{b(1+b)+a^2}{a(1+b)}\]

OpenStudy (anonymous):

now some more algebra, but just a little \[\frac{b(1+b)+a^2}{a(1+b)}=\frac{b+b^2+a^2}{a(1+b)}\]

OpenStudy (anonymous):

that is just the distributive law in the numerator now that we are done with algebra, we can go back to sines and cosines

OpenStudy (anonymous):

sooooooo then \[\frac{ \cos \theta }{ \sin \theta } + \frac{ \cos \theta }{ 1+\sin \theta }\] should become....

OpenStudy (anonymous):

\[\frac{\sin(x)+\sin^2(x)+\cos^2(x)}{\cos(x)(1+\sin(x))}\]

OpenStudy (anonymous):

Ok. But now comes an even trickier part.... I have to find a way to reduce that

OpenStudy (anonymous):

hold a sec the initial question was \[\frac{ \sin \theta }{ \cos \theta }+\frac{ \cos \theta }{ 1+ \sin \theta }\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so after the algebra we get to \[\frac{\sin(x)+\sin^2(x)+\cos^2(x)}{\cos(x)(1+\sin(x))}\] don't look to reduce yet

OpenStudy (anonymous):

the usual miracle occurs, \[\sin^2(x)+\cos^2(x)=1\] the mother of all trig identities

OpenStudy (anonymous):

correct

OpenStudy (chrisplusian):

The reasoning behind adding and subtracting the fractions is exactly what was said above, but the step being left out from above that might make it clear is that you just multiply both fractions by a number that will make both denominators the same or "common". In the above example....\[\frac{ b }{ a }+\frac{ a }{ 1+b } = (\frac{ 1+b }{ 1+b })(\frac{ b }{ a })+(\frac{ a }{ a })(\frac{ a }{ 1+b })\] and that is where the other equation @satellite73 came up with came from. Not to inturrupt

OpenStudy (anonymous):

\[\frac{\sin(x)+\overbrace{\sin^2(x)+\cos^2(x)}^{\text{ this is 1}}}{\cos(x)(1+\sin(x))}\]

OpenStudy (anonymous):

\[\frac{\sin(x)+1}{\cos(x)(1+\sin(x))}\] NOW you can cancel the common factor of \[1+\sin(x)\] top and bottom

OpenStudy (anonymous):

Ok, ok, I see it now. so the Sins actually match and leave you with \[\frac{ 1 }{ \cos \theta }\]

OpenStudy (anonymous):

yes if you want to impress your teacher, write it as \[\sec(\theta)\]

OpenStudy (anonymous):

i hope you also see that 98% of this is algebra

OpenStudy (anonymous):

which, if your algebra is not what it needs to be, is going to be a problem, so bone up on it also use the gimmick of replacing sine and cosine by letters to make the algebra easier if we had to do this writing sine and cosine each time it would have been a pain